An achievement test was administer to a class of 20,000 students. The mean score was 80 and the standard deviation was 11. If Lingard scored 72 in the test, how many students did better than him?
1
Expert's answer
2021-07-19T05:49:04-0400

"N=20000 \\

\mu=80 \\

\sigma=11 \\

P(X>72) = 1 -P(X<72) \\

= 1 -P(Z< \frac{72-80}{11}) \\

= 1 -P(Z < -0.7272) \\

= 1 -0.2336 \\

= 0.7664 \\

n(X>72) = 20000 \times 0.7664 = 15328"

Answer: 15328 students were better than Lingard